OpenStudy (anonymous):

How to find the exact value of sin50 cos40 + sin40cos50?

OpenStudy (mathstudent55):

\(\sin (\alpha + \beta) = \sin \alpha \cos \beta + \sin \beta \cos \alpha \)

OpenStudy (mathstudent55):

That is the angle addition formula for sine.

OpenStudy (anonymous):

Please, elaborate it. :(

OpenStudy (mathstudent55):

Use the formula above and replace every alpha with 50 and every beta with 40. What do you get?

OpenStudy (anonymous):

sin90

OpenStudy (mathstudent55):

\(\sin \color{red}{(40 + 50)} = \sin 50 \cos 40 + \sin 40 \cos 50\) \(\sin \color{red}{90} = \sin 50 \cos 40 + \sin 40 \cos 50\) You do know an exact value for \(\sin 90^o\), right?

OpenStudy (anonymous):

Its 1.

OpenStudy (mathstudent55):

Correct.